OpenStudy (anonymous):

An experiment consists of rolling two fair(not weighted) dice and adding the dots on the two sides facing up. Each die has the number 1 on two opposite faces, the number 2 on two opposite faces, and the number 3 on twoopposite faces. Compute the probability of obtaining the indicated sums in problem. 41) 4

OpenStudy (anonymous):

adding the dots on the two sides facing up? is that how those things work?

OpenStudy (anonymous):

there are a total of 36 possible outcomes, just like with regular dice. if you count, you will see that 12 of them give a total of 4, so your answer is \[\frac{12}{36}=\frac{1}{3}\]

OpenStudy (anonymous):

the rest can also be obtained by counting. the probability you get a sum of 2 is \[\frac{4}{36}\] likewise for a sum of 6 and the probability you get a sum of 3 and the probability you get a sum of 5 is \[\frac{8}{36}\]